OpenStudy (anonymous):

b) The number of ants in an colony increases exponentially at a rate of 4.6% per week. How long do you have to wait until the ant colony has tripled in size?

OpenStudy (atlas):

x(1+0.046)^n = 3x

OpenStudy (atlas):

FInd n

OpenStudy (anonymous):

gimme a sec

OpenStudy (atlas):

The concept is after 1 week, the size of ant colony will be x + 4.6% of x = x(1+0.046)

OpenStudy (atlas):

Let x(1+1.046) = y In the second week again it will become y + 4.6% of y = y(1+0.046) = x(1+0.046)(1+0.046) = x(1+0.046)^2

OpenStudy (atlas):

and so on

OpenStudy (campbell_st):

well I'd use \[P_{n} = P_{0} \times e^kt\] and for the population to triple \[P_{0} = 1 ... and..... P_{n} = 3\] so the equation becomes \[3 = 4^{0.046 \times t}\] hopefully that helps you solve for the time...

OpenStudy (atlas):

After n weeks -> x(1+0.046)^n which should be equal to 3x

OpenStudy (campbell_st):

oops the equation should read \[3 = e^{0.046\times t}\] solve for t

OpenStudy (atlas):

@cambell_st : I don't think exponential growth means what you think http://en.wikipedia.org/wiki/Exponential_growth

OpenStudy (atlas):

@campbell_st

OpenStudy (atlas):

not to be confused with e being called an exponential function

OpenStudy (campbell_st):

well if you use the phrase exponential growth you need a base and exponent... I found it would take approx 24 weeks and using the compound interest formula as you have done... which is also a growth model... the time is 24.4.... so rounding.... or approx 24 weeks... ummm... which is correct...

OpenStudy (campbell_st):

and here is a site that talks about growth and decay in terms of e http://www.mathsisfun.com/algebra/exponential-growth.html

OpenStudy (campbell_st):

the only determining factor is whether you are told that the rate of growth is proportional to itself... so who knows

OpenStudy (atlas):

i get your viewpoint now i used the above form of exponential growth because the growth is mentioned in percentage

OpenStudy (atlas):

*growth rate